# FreeScale: Unleashing the Resolution of Diffusion Models via Tuning-Free Scale Fusion

### v1

@@ Abstract
Visual diffusion models achieve remarkable progress, yet they are typically trained at limited resolutions due to the lack of high-resolution data and constrained computation resources, hampering their ability to generate high-fidelity images or videos at higher resolutions. Recent efforts have explored tuning-free strategies to exhibit the untapped potential higher-resolution visual generation of pre-trained models. However, these methods are still prone to producing low-quality visual content with repetitive patterns. The key obstacle lies in the inevitable increase in high-frequency information when the model generates visual content exceeding its training resolution, leading to undesirable repetitive patterns deriving from the accumulated errors. To tackle this challenge, we propose FreeScale, a tuning-free inference paradigm to enable higher-resolution visual generation via scale fusion. Specifically, FreeScale processes information from different receptive scales and then fuses it by extracting desired frequency components. Extensive experiments validate the superiority of our paradigm in extending the capabilities of higher-resolution visual generation for both image and video models. Notably, compared with previous best-performing methods, FreeScale unlocks the generation of 8k-resolution images for the first time.
Figure 1: Gallery of FreeScale. Original SDXL citesdxl can only generate images with a resolution of up to $1024^{2}$ without losing quality, while FreeScale successfully extends SDXL to generate $8192^{2}$ images without any fine-tuning. All generated images are produced using a single A800 GPU. Best viewed ZOOMED-IN.
11footnotetext: Project Leader22footnotetext: Corresponding Authors
@@ 1 Introduction
  • We unlock the generation of 8k-resolution images for the first time. Compared to other state-of-the-art tuning-free methods, FreeScale gains superior visual quality with less inference time.
@@ 2 Related Work
Higher-Resolution Visual Generation. High-resolution visual synthesis is classic challenging in the generative field due to the difficulty of collecting plenty of high-resolution data and the requirement of substantial computational resources. Recent methods for higher-resolution generation can mainly be divided into two categories: 1) training/tuning methods with high-resolution data and large models [43, 22, 38, 33, 14, 51], or 2) tuning-free methods without any additional data requirement [18, 13, 16, 31, 30, 27, 25, 7]. Training with high-resolution data on larger models should be a more fundamental solution. However, high-resolution visual data, especially high-resolution videos, only accounts for a small proportion. Meanwhile, targeting for modeling higher-resolution data demands a notably increased requirement in model capacity. Based on current data and calculation resources, tuning-free approaches are more achievable for high-resolution generation.
One straightforward approach is to generate video patches of the same resolution as the training video and then stitch them together. Although eliminating the training-inference gap, this method results in disconnected and incoherent patches. MultiDiffusion [2] addresses this issue by smoothly fusing patches during the denoising process. DemoFusion [13] utilizes this mechanism and adds global perception to ensure the rationality of the overall layout. However, this solution easily leads to small object repetition generation. ScaleCrafter [18] argues that the object repetition issue is mainly caused by the limited convolutional receptive field and uses dilated convolutional layers to enlarge the convolutional receptive field. Although successfully removing small object repetition, ScaleCrafter suffers from a new problem of local repetition. FouriScale [23] concludes that all kinds of repetitions are from the nonalignment of frequency domain in different scales. FouriScale removes the high-frequency signals of the latent before the convolution operation and achieves no repetition at all. But this violent editing operation on the frequency domain leads to weird results with unnatural colors and textures. Another solution is directly removing the text semantics from unexpected areas in the input level [32, 34]. However, it only works for small object repetition and will suffer information leakage through the temporal layers in the video generation.
@@ 3.1 Preliminaries
Latent Diffusion Models (LDM) first encodes an given image ${\mathbf{x}}$ to the latent space ${\mathbf{z}}$ via the encoder of the pre-trained auto-encoder $\mathcal{E}$: $z=\mathcal{E}(x)$. Then a forward diffusion process is used to gradually add noise to the latent data ${\mathbf{z}}_{0}\sim p({\mathbf{z}}_{0})$ and learn a denoising model to reverse this process. The forward process contains $T$ timesteps, which gradually add noise to the latent sample ${\mathbf{z}}_{0}$ to yield ${\mathbf{z}}_{t}$ through a parameterization trick:
@@ 3.2 Tailored Self-Cascade Upscaling
where $\tilde{{\mathbf{z}}}$ means the noised intermediate latent, $r$ is the resolution level ($1$ represents original resolution, $2$ represents the twice height and width), and $\phi$ is an upsampling operation. In this way, the framework will generate a reasonable visual structure in low resolution and maintain the structure when generating higher-resolution results.
@@ 3.3 Restrained Dilated Convolution
ScaleCrafter [18] observes that the primary cause of the object repetition issue is the limited convolutional receptive field and proposes dilated convolution to solve it. Given a hidden feature map ${\mathbf{h}}$, a convolutional kernel $\bm{k}$, and the the dilation operation $\Phi_{d}(\cdot)$ with factor $d$, the dilated convolution can be represented as:
@@
FID $\downarrow$ KID $\downarrow$ $\text{FID}_{c}$downarrow $\text{KID}_{c}$downarrow Time (min) $\downarrow$ FID $\downarrow$ KID $\downarrow$ $\text{FID}_{c}$downarrow $\text{KID}_{c}$downarrow Time (min) $\downarrow$
SDXL-DI [36] 64.313 0.008 31.042 0.004 0.648 134.075 0.044 42.383 0.009 5.456
ScaleCrafter [18] 67.545 0.013 60.151 0.020 0.653 100.419 0.033 116.179 0.053 9.255
DemoFusion [13] 65.864 0.016 63.001 0.024 1.441 72.378 0.020 94.975 0.045 11.382
FouriScale [23] 68.965 0.016 69.655 0.026 1.224 93.079 0.029 128.862 0.068 8.446
Ours 44.723 0.001 36.276 0.006 0.853 49.796 0.004 71.369 0.029 6.240
@@ 4 Experiments
Figure 4: Qualitative results of flexible control for detail level. A better result will be generated by adding the coefficient weight in the area of Griffons and reducing the coefficient weight in the other regions. Best viewed ZOOMED-IN.
@@
Experimental Settings. We conduct experiments based on an open-source T2I diffusion model SDXL [36] and an open-source T2V diffusion model VideoCrafter2 [10]. Considering the computing resources that can be afforded, we evaluate the image generation at resolutions of $2048^{2}$ and $4096^{2}$, and video generation at resolutions of $640$. All experiments are produced using a single A800 GPU.
Evaluation Metrics. Since higher-resolution inference methods are intended to maintain the quality of the original resolution outputs, we calculate all metrics between the originally generated low-resolution images/videos and the corresponding high-resolution outputs. To evaluate the quality of generated images, we report Frechet Image Distance (FID) [19], and Kernel Image Distance (KID)  [4]. FID and KID need to resize the images to $299$ before the comparison and this operation may cause quality loss for high-resolution images. Inspired by previous work [8], we also use cropped local patches to calculate these metrics without resizing, termed $\text{FID}_{c}$ and $\text{KID}_{c}$. We use Frechet Video Distance (FVD) [44] to evaluate the quality of video generation. In addition, we test dynamic degree and aesthetic quality from the VBench [24] to evaluate the dynamics and aesthetics.
Method $2048^{2}$ $4096^{2}$
FID $\downarrow$ KID $\downarrow$ $\text{FID}_{c}$downarrow $\text{KID}_{c}$downarrow Time (min) $\downarrow$ FID $\downarrow$ KID $\downarrow$ $\text{FID}_{c}$downarrow $\text{KID}_{c}$downarrow Time (min) $\downarrow$
w/o Scale Fusion 75.717 0.017 76.536 0.026 0.614 68.115 0.012 100.065 0.037 4.566
Dilated Up-Blocks 75.372 0.017 76.673 0.025 0.861 67.447 0.011 98.558 0.035 6.245
Latent Space Upsampling 72.454 0.015 71.793 0.023 0.840 65.081 0.009 88.632 0.029 6.113
Ours 44.723 0.001 36.276 0.006 0.853 49.796 0.004 71.369 0.029 6.240
@@ 4.1 Higher-Resolution Image Generation
The quantitative results also confirm the superiority of FreeScale. As shown in Table 1, SDXL-DI achieves the best $\text{FID}_{c}$ and $\text{KID}_{c}$. The reason is that SDXL-DI tends to generate multiple duplicated objects and its crop may be closer to the reference images. However, this behavior will sacrifice the visual structure thus SDXL gains the worst FID and KID in the resolution of $4096^{2}$. Overall, our approach achieves the best or second-best scores for all quality-related metrics with negligible additional time costs.
In addition, FreeScale provides flexible control for detail level in generated results. Figure 4 shows a demo of changing the detail level of different semantic areas. After we get the upsampled 1×1$\times$1 × results, it is easy to calculate semantic masks [28] and assign different $\alpha$ for each region in Equation 4. As shown in Figure 4, we will obtain a better result when we add the coefficient weight in the area of Griffons and reduce the coefficient weight in other regions.
@@
Figure 6: Qualitative image comparisons with ablations. Our full method performs the best. The resolution of results is $4096^{2}$ for better visualizing the difference between the various strategies. Best viewed ZOOMED-IN.
@@ 4.2 Higher-Resolution Video Generation
As shown in Figure 5, the behavior of VC2-DI and ScaleCrafter are similar to the corresponding version in image generation, tending to generate duplicated whole objects and local parts, respectively. However, DemoFusion has completely unexpected behavior in the video generation. Its Dilated Sampling mechanism brings strange patterns all over the frames and Skip Residual operation makes the whole video blur. In contrast, our FreeScale effectively generates higher-resolution videos with high fidelity. Table 3 exhibits that our method achieves the best FVD with a shorter inference time. More quantitative evaluations can be found in the supplementary file.
@@ 5 Conclusion
This study introduces FreeScale, a tuning-free inference paradigm designed to enhance high-resolution generation capabilities in pre-trained diffusion models. By leveraging multi-scale fusion and selective frequency extraction, FreeScale effectively addresses common issues in high-resolution generation, such as repetitive patterns and quality degradation. Experimental results demonstrate the superiority of FreeScale in both image and video generation, surpassing existing methods in visual quality while also having significant advantages in inference time. Compared to prior approaches, FreeScale not only eliminates various forms of visual repetition but also ensures detail clarity and structural coherence in generated visuals. Eventually, FreeScale achieves unprecedented 8k-resolution image generation.
@@ 6 Acknowledgements
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2022-01-035T), the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221- 0012), NTU NAP, and Alibaba Group. Special thanks to Yingqing He and Lanqing Guo for sharing their invaluable experience and advice in getting us started quickly in the higher-resolution visual generation task.
@@
Overview. In the supplementary material, we introduce implementation details in Section A, show more evaluation results in Section B, exhibit gallery of 8k images in Section C, and finally, discuss limitation and future work in Section D.
@@ Appendix A Implementation Details
For tailored self-cascade upscaling, we set $K=700$ in Equation 3 for all experiments. And in Equation 4, $\alpha$ is set as a scaler, $2$, by default. To avoid excessive and messy textures in generating 8k images, $\alpha$ is reduced to $1$. In Figure 4, $\alpha$ is $3$ and $0.5$ in the targeted and other areas, respectively.
For restrained dilated convolution, the dilation factor $d$ in Equation 5 is equal to the resolution level ($1$ represents original resolution, $2$ represents the twice height and width).
@@ Appendix B More Evaluation
B.1 Ablation Study for Video Generation
B.2 User Study
In addition, we conducted a user study to evaluate our results on human subjective perception. Users are asked to watch the generated images of all the methods, where each example is displayed in a random order to avoid bias, and then pick up the best one in three evaluation aspects. A total of $23$ users were asked to pick the best one according to the image-text alignment, image quality, and visual structure, respectively. As shown in Table  5, our approach gains the most votes for all aspects, outperforming baseline methods by a large margin.
Dilated Up-Blocks 523.323 0.363 0.611 3.788
RGB Upsampling 422.245 0.381 0.604 3.799
Ours 484.711 0.383 0.621 3.787
@@ B.2 User Study
Appendix C Gallery of 8k Images

### v2

@@ Abstract
Visual diffusion models achieve remarkable progress, yet they are typically trained at limited resolutions due to the lack of high-resolution data and constrained computation resources, hampering their ability to generate high-fidelity images or videos at higher resolutions. Recent efforts have explored tuning-free strategies to exhibit the untapped potential higher-resolution visual generation of pre-trained models. However, these methods are still prone to producing low-quality visual content with repetitive patterns. The key obstacle lies in the inevitable increase in high-frequency information when the model generates visual content exceeding its training resolution, leading to undesirable repetitive patterns deriving from the accumulated errors. To tackle this challenge, we propose FreeScale, a tuning-free inference paradigm to enable higher-resolution visual generation via scale fusion. Specifically, FreeScale processes information from different receptive scales and then fuses it by extracting desired frequency components. Extensive experiments validate the superiority of our paradigm in extending the capabilities of higher-resolution visual generation for both image and video models. Notably, compared with previous best-performing methods, FreeScale unlocks the 8k-resolution text-to-image generation for the first time.
Figure 1: Gallery of FreeScale. Original SDXL [40] can only generate images with a resolution of up to $1024^{2}$ without losing quality, while FreeScale successfully extends SDXL to generate $8192^{2}$ images without any fine-tuning. All generated images are produced using a single A800 GPU. Best viewed ZOOMED-IN.
@@ 1 Introduction
  • Compared to other state-of-the-art tuning-free methods, we unlock the 8k-resolution text-to-image generation for the first time.
@@ 2 Related Work
Higher-Resolution Visual Generation. High-resolution visual synthesis is a classic challenge in the generative field due to the difficulty of collecting plenty of high-resolution data and the requirement of substantial computational resources. Recent methods for higher-resolution generation can mainly be divided into two categories: 1) training/tuning methods with high-resolution data and large models [49, 24, 42, 37, 16, 59, 12], or 2) tuning-free methods without any additional data requirement [18, 35, 34, 29, 27, 7, 58, 31]. Training with high-resolution data on larger models should be a more fundamental solution. However, high-resolution visual data only accounts for a small proportion. Meanwhile, targeting for modeling higher-resolution data demands a notably increased requirement in model capacity. Based on current data and calculation resources, tuning-free approaches are more achievable for high-resolution generation.
One straightforward approach is to generate visual patches of the same resolution as the training data and then stitch them together. Although eliminating the training-inference gap, this method results in disconnected and incoherent patches. MultiDiffusion [2] addresses this problem by fusing patches smoothly during the denoising process. DemoFusion [14] utilizes this mechanism and adds global perception to ensure the rationality of the overall layout. However, this solution easily leads to the generation of small object repetition. ScaleCrafter [20] argues that the object repetition issue is mainly caused by the limited convolutional receptive field and uses dilated convolutional layers to enlarge the convolutional receptive field. Although successful in removing small object repetition, ScaleCrafter suffers from a new problem of local repetition. FouriScale [25] concludes that all types of repetitions are from the non-alignment of frequency domain on different scales. FouriScale removes the high-frequency signals of the latent prior to convolution operation and achieves no repetition at all. But this violent editing operation on the frequency domain leads to strange results with unnatural colors and textures. Another solution is directly removing the text semantics from unexpected areas in the input level [36, 38]. However, it only works for small object repetition and will suffer information leakage through the temporal layers in the video generation. With the additional pose as input, BeyondScene [30] has achieved 8k human image generation. However, its scope is limited to human image generation due to the requirement of additional pose input. FreeScale is the first 8k-resolution text-to-image generation method without these constraints.
@@ 3.1 Preliminaries
Latent Diffusion Models (LDM) first encodes a given image ${\mathbf{x}}$ to the latent space ${\mathbf{z}}$ via the encoder of the pre-trained auto-encoder $\mathcal{E}$: $z=\mathcal{E}(x)$. Then a forward diffusion process is used to gradually add noise to the latent data ${\mathbf{z}}_{0}\sim p({\mathbf{z}}_{0})$ and learn a denoising model to reverse this process. The forward process contains $T$ timesteps, which gradually add noise to the latent sample ${\mathbf{z}}_{0}$ to yield ${\mathbf{z}}_{t}$ through a parameterization trick:
@@ 3.2 Tailored Self-Cascade Upscaling
where $\tilde{{\mathbf{z}}}$ means the noised intermediate latent, $r$ is the resolution level ($1$ represents original resolution, $2$ represents the twice height and width), and $\phi$ is an upsampling operation. Specifically, FreeScale will denoise using the training resolution. The intermediate results will then be gradually up-sampled. In the higher resolution, blurry details from the upsampling will be removed by adding noise (to the level of timestep $K$) and denoising. In this way, the framework will generate a reasonable visual structure in low resolution and maintain the structure when generating higher-resolution results.
@@ 3.3 Restrained Dilated Convolution
ScaleCrafter [20] observes that the primary cause of the object repetition issue is the limited convolutional receptive field and proposes dilated convolution to solve it. Given a hidden feature map ${\mathbf{h}}$, a convolutional kernel $\bm{k}$, and the dilation operation $\Phi_{d}(\cdot)$ with factor $d$, the dilated convolution can be represented as:
@@
FID $\downarrow$ KID $\downarrow$ $\text{FID}_{c}$downarrow $\text{KID}_{c}$downarrow IS $\uparrow$ Time (min) $\downarrow$ FID $\downarrow$ KID $\downarrow$ $\text{FID}_{c}$downarrow $\text{KID}_{c}$downarrow IS $\uparrow$ Time (min) $\downarrow$
SDXL-DI [40] 64.313 0.008 31.042 0.004 10.424 0.648 134.075 0.044 42.383 0.009 7.036 5.456
ScaleCrafter [20] 67.545 0.013 60.151 0.020 11.399 0.653 100.419 0.033 116.179 0.053 8.805 9.255
DemoFusion [14] 65.864 0.016 63.001 0.024 13.282 1.441 72.378 0.020 94.975 0.045 12.450 11.382
FouriScale [25] 68.965 0.016 69.655 0.026 11.055 1.224 93.079 0.029 128.862 0.068 8.248 8.446
Ours 44.723 0.001 36.276 0.006 12.747 0.853 49.796 0.004 71.369 0.029 12.572 6.240
@@ 4 Experiments
Figure 4: Results of flexible control for detail level. A better result will be generated by adding the coefficient weight in the area of Griffons and reducing the coefficient weight in the other regions. Best viewed ZOOMED-IN.
Figure 5: Results of local semantic editing. FreeScale makes the hair purple or edits the face to make this person look more Japanese in the higher-resolution ($4096^{2}$).
Figure 6: Qualitative image comparisons with ablations. Our full method performs the best. The resolution of results is $4096^{2}$ for better visualizing the difference between the various strategies.
@@
Experimental Settings. We conduct experiments based on an open-source T2I diffusion model SDXL [40] and an open-source T2V diffusion model VideoCrafter2 [10]. Considering the computing resources that can be afforded, we evaluate the image generation at resolutions of $2048^{2}$ and $4096^{2}$, and video generation at resolutions of $640\times 1024$. All experiments are produced using a single A800 GPU.
Evaluation Metrics. Since higher-resolution inference methods are intended to maintain the quality of the original resolution outputs, we calculate all metrics between the originally generated low-resolution images/videos and the corresponding high-resolution outputs. To evaluate the quality of generated images, we report Frechet Image Distance (FID) [21], Kernel Image Distance (KID)  [4] and IS (Inception Score) [44]. FID and KID need to resize the images to $299$ before the comparison and this operation may cause quality loss for high-resolution images. Inspired by previous work [8], we also use cropped local patches to calculate these metrics without resizing, termed $\text{FID}_{c}$ and $\text{KID}_{c}$. We use Frechet Video Distance (FVD) [50] to evaluate the quality of video generation. In addition, we test dynamic degree and aesthetic quality from the VBench [26] to evaluate the dynamics and aesthetics.
@@ 4.1 Higher-Resolution Image Generation
The quantitative results also confirm the superiority of FreeScale. As shown in Table 1, SDXL-DI achieves the best $\text{FID}_{c}$ and $\text{KID}_{c}$. The reason is that SDXL-DI tends to generate multiple duplicated objects and its crop may be closer to the reference images. However, this behavior will sacrifice the visual structure thus SDXL gains the worst FID, KID and IS in the resolution of $4096^{2}$. Overall, our approach achieves the best or second-best scores for all quality-related metrics with negligible additional time costs.
Local Control. FreeScale provides flexible control for detail level in generated results. Figure 4 shows a demo of changing the detail level of different semantic areas. During the process of tailored self-cascade upscaling, we will get 1×1$\times$1 × results as intermediates. Although more details will be added or modified in the later higher-resolution stages, the overall structure and main content of the image have been determined in the 1×1$\times$1 × results. It is easy to calculate semantic masks [32] and assign different $\alpha$ for each region in Equation 4. As shown in Figure 4, we will obtain a better result when we add the coefficient weight in the area of Griffons and reduce the coefficient weight in other regions.
In addition, this mechanism can even be extended to local semantic editing. Utilizing semantic mask from 1×1$\times$1 × results, we can inject different text semantics into different regions in the layers of cross-attention. As shown in Figure 5, FreeScale successfully edits the hair and face in the higher-resolution results.
@@
Method $2048^{2}$ $4096^{2}$
FID $\downarrow$ KID $\downarrow$ $\text{FID}_{c}$downarrow $\text{KID}_{c}$downarrow IS $\uparrow$ Time (min) $\downarrow$ FID $\downarrow$ KID $\downarrow$ $\text{FID}_{c}$downarrow $\text{KID}_{c}$downarrow IS $\uparrow$ Time (min) $\downarrow$
w/o Scale Fusion 75.717 0.017 76.536 0.026 12.743 0.614 68.115 0.012 100.065 0.037 12.415 4.566
Dilated Up-Blocks 75.372 0.017 76.673 0.025 12.541 0.861 67.447 0.011 98.558 0.035 12.543 6.245
Latent Space Upsampling 72.454 0.015 71.793 0.023 12.210 0.840 65.081 0.009 88.632 0.029 11.307 6.113
Ours 44.723 0.001 36.276 0.006 12.747 0.853 49.796 0.004 71.369 0.029 12.572 6.240
@@ 4.2 Higher-Resolution Video Generation
As shown in Figure 7, the behavior of VC2-DI and ScaleCrafter are similar to the corresponding version in image generation, tending to generate duplicated whole objects and local parts, respectively. However, DemoFusion has completely unexpected behavior in the video generation. Its Dilated Sampling mechanism brings strange patterns all over the frames and Skip Residual operation makes the whole video blur. In contrast, our FreeScale effectively generates higher-resolution videos with high fidelity. Table 2 exhibits that our method achieves the best FVD, dynamic degree and aesthetic quality. In addition, the time cost saved by skipping certain timesteps near pure noise (transparent blocks in Figure 2) even outweighs the extra time caused by other modules in FreeScale.
@@ 5 Conclusion
This study introduces FreeScale, a tuning-free inference paradigm designed to enhance high-resolution generation capabilities in pre-trained diffusion models. By leveraging multi-scale fusion and selective frequency extraction, FreeScale effectively addresses common issues in high-resolution generation, such as repetitive patterns and quality degradation. Experimental results demonstrate the superiority of FreeScale in both image and video generation, surpassing existing methods in visual quality while also having significant advantages in inference time. FreeScale not only eliminates various forms of visual repetition but also ensures detail clarity and structural coherence in generated visuals. Additional local control capabilities provide users with more flexibility. Eventually, FreeScale achieves unprecedented 8k-resolution text-to-image generation.
@@ 6 Acknowledgements
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2022-01-035T), the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221-0012, MOE-T2EP20223-0002), and Alibaba Group. Special thanks to Yingqing He and Lanqing Guo for sharing their invaluable experience and advice in getting us started quickly in the higher-resolution visual generation task.
@@
Overview. In the supplementary material, we introduce implementation details in Section A, show more evaluations in Section B, exhibit more results in Section C, and finally, discuss limitations and future work in Section D.
@@ Appendix A Implementation Details
For tailored self-cascade upscaling, we set $K=700$ in Equation 3 for all experiments. And in Equation 4, $\alpha$ is set as a scaler, $2$, by default. To avoid excessive and messy textures in generating 8k images, $\alpha$ is reduced to $1$. In Figure 4, $\alpha$ is $3$ and $0.5$ in the targeted and other areas, respectively. Users can further adjust these parameters according to the detailed requirements of different images. For restrained dilated convolution, the dilation factor $d$ in Equation 5 is equal to the resolution level ($1$ represents original resolution, $2$ represents the twice height and width). For scale fusion, the kernel size is $12\times\sqrt{height\times width\div(1024\times 1024)}-$ 1 and the standard deviation is $1$ in Equation 7.
@@ Appendix B More Evaluation
Table 4: Image quantitative comparisons with super-resolution. Compared to super-resolution post-processing setting SDXL+Real-ESRGAN, FreeScale also achieves competitive performance. As reported in most previously published related works, higher-resolution generation methods are hard to beat SR methods completely on quantitative metrics due to the difference in difficulty between the two tasks.
Method FID $\downarrow$ KID $\downarrow$ $\text{FID}_{c}$downarrow $\text{KID}_{c}$downarrow IS $\uparrow$
SDXL+Real-ESRGAN [52] 43.476 0.000 73.524 0.024 12.599
Ours 49.796 0.004 71.369 0.029 12.572
@@
Table 6: User study for Video Generation. Users are required to pick the best one among our proposed FreeScale with the other baseline methods in terms of text alignment, cover quality, and video quality.
Method Text Alignment Cover Quality Video Quality
VC2-DI 5.38% 4.62% 3.85%
ScaleCrafter 4.62% 5.38% 0.77%
DemoFusion 30.00% 26.92% 30.77%
Ours 60.00% 63.08% 64.62%
B.1 Comparison with Super-Resolution
Different from traditional super-resolution (SR) tasks. Higher-resolution generation aims to tap the potential of the pre-trained model itself. Therefore, the performance of the higher-resolution generation method is based on the base model rather than another additional SR model. We compare our method with a super-resolution post-processing setting: SDXL+Real-ESRGAN [52]. As shown in Table 4, FreeScale achieves competitive performance in quantitative metrics. As reported in most previously published related works [20, 14], higher-resolution generation methods are hard to beat SR methods completely on quantitative metrics due to the difference in difficulty between the two tasks. However, Figure 8 shows that FreeScale is not inferior to SDXL+Real-ESRGAN in visual quality, and adds more details. In addition, SR methods will faithfully follow the low-resolution input while FreeScale can regenerate the original blurred areas based on the prior knowledge that the model has learned (the eyes and logos in Figure 8).
Figure 8: Image qualitative comparisons with super-resolution. FreeScale is not inferior to SDXL+Real-ESRGAN in visual quality, and adds more details. In addition, SR methods will faithfully follow the low-resolution input while FreeScale can regenerate the original blurred areas based on the prior knowledge that the model has learned. Best viewed ZOOMED-IN.
B.2 User Study
In addition, we conducted a user study to evaluate our results on human subjective perception. Users are asked to watch the generated images of all the methods, where each example is displayed in a random order to avoid bias, and then pick the best one in three evaluation aspects. A total of $23$ users were asked to pick the best one according to the image-text alignment, image quality, and visual structure, respectively. As shown in Table 5, our approach gains the most votes for all aspects, outperforming baseline methods by a large margin.
We also add a human study for video generation. Users were asked to pick the best one according to the text alignment, cover quality, and video quality, respectively. As shown in Table 6, our method still gains the most votes for all aspects, outperforming baseline approaches significantly.
Dilated Up-Blocks 523.323 0.363 0.611 3.788
RGB Upsampling 422.245 0.381 0.604 3.799
Ours 484.711 0.383 0.621 3.787
B.3 Ablation Study for Video Generation
Appendix C More Results
Figure 9: Flexible aspect ratio generation. FreeScale can directly achieve a flexible aspect ratio (the resolution must be a multiple of $512$) without any adaptation.
Figure 11: Fast generation with SDXL-Turbo. FreeScale can help SDXL-Turbo generate results at $2048^{2}$ resolution with even $2$ timesteps.
Figure 13: Structure gap. UNet-based LDMs and DiT-based LDMs will face different challenges in the higher-resolution generation task. UNet-based LDMs face repetition problems while DiT-based LDMs face blur problems.
Table 8: Image quantitative comparisons with baselines in $2048\times 4096$ resolution. FreeScale still achieves the best or second-best scores for all metrics.
Method FID $\downarrow$ KID $\downarrow$ $\text{FID}_{c}$downarrow $\text{KID}_{c}$downarrow IS $\uparrow$
SDXL-DI 97.493 0.026 38.273 0.009 7.258
ScaleCrafter 97.235 0.032 107.582 0.050 8.001
DemoFusion 72.196 0.019 91.264 0.044 10.622
FouriScale 95.891 0.032 118.306 0.061 8.422
Ours 54.704 0.004 65.584 0.025 11.323
C.1 Flexible Aspect Ratio Generation
As shown in Figure 9, FreeScale can directly achieve a flexible aspect ratio (the resolution must be a multiple of $512$) without any adaptation. We also add quantitative experiments for $2048\times 4096$ resolution. As shown in Table 8, FreeScale still achieves the best or second-best scores for all metrics.
C.2 Fast Generation with SDXL-Turbo
FreeScale can easily be compatible with other models with similar structures. SDXL-Turbo [45] is a distilled version of SDXL [40] and can produce similar quality results with $2\sim 4$ timesteps. However, SDXL-Turbo can only generate results at $512^{2}$ resolution due to the knowledge loss during distillation. As shown in Figure 11, FreeScale can help SDXL-Turbo generate results at $2048^{2}$ resolution.
C.3 Gallery of 8k Images